A solenoid with 500·mA of current produces a magnetic field of 100·gauss. If you change the current to 200·mA, the magnetic field will be gauss.
Answer #1

Megnetic field due to solenoid =onI

n-number of turns per unit length

o- constant

So if you decrease the current to 200/500=2/5 factor

Megnetic field will be (2/5)*100=40 Gauss